Whirly Corporation’s contribution format income statement for the most recent month is shown below: Total Per Unit Sales (7,600 units) $ 235,600 $ 31.00 Variable expenses 152,000 20.00 Contribution margin 83,600 $ 11.00 Fixed expenses 54,400 Net operating income $ 29,200 Required: (Consider each case independently): 1. What would be the revised net operating income per month if the sales volume increases by 90 units? 2. What would be the revised net operating income per month if the sales volume decreases by 90 units? 3. What would be the revised net operating income per month if the sales volume is 6,600 units?
Answer #1
Case 1 Case 2 Case 3
Unit Sales (7600+90)=7690 (7600-90)=7510 6600
Sales (7690*31)=238390 (7510*31)=232810 (6600*31)=204600
Variable expenses (7690*20)=153800 (7510*20)=150200 (6600*20)=132,000
Contribution margin 84590 82610 72600
Fixed expenses 54400 54400 54400
Revised net operating income $30190 $28210 $18200